A train at 72km/h undergoes a uniform retardation of 2m/s when brakes are applied. Find the distance traveled from the place where the brakes were applied.
1
Expert's answer
2019-04-18T10:06:50-0400

Let's first convert km/h to m/s:


"v_0 = 72 \dfrac{km}{h} \cdot \dfrac{1000 m}{1 km} \cdot \dfrac{1 h}{3600 s} = 20 \dfrac{m}{s}."

We can find the distance traveled by the train from the place where the brakes were applied from the kinematic equation:


"v_f^2 = v_0^2 + 2ad,"

here, "v_0 = 20 m\/s" is the initial velocity of the train, "v_f = 0" is the final velocity of the train, "a = -2 m\/s^2" is the deceleration of the train, "d" is the distance traveled by the train from the place where the brakes were applied.

Then, we get:


"d = \dfrac{v_f^2 - v_0^2}{2a},""d = \dfrac{(0 \dfrac{m}{s})^2 - (20 \dfrac{m}{s})^2}{2 \cdot (-2 \dfrac{m}{s^2})} = 100 m."

Answer:

"d = 100 m."